OpenStudy (anonymous):

need help with -9p-15=93 ??

OpenStudy (anonymous):

add 15 to both sides

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

you got it from there ??

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

93 + 15 = ???

OpenStudy (anonymous):

-9p-15=93 -15 -15 p=78 ?

OpenStudy (anonymous):

nooo you add 15 to both sides

OpenStudy (anonymous):

so like 9p-5=93 =1 5

OpenStudy (anonymous):

-9p=108

OpenStudy (anonymous):

now divide -9 from both sides